### v1

@@ 6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.39 admits a unique adapted solution. Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.37 guarantees the existence and uniqueness of solution. Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.39) with the same formulation and the same $(\rm H_{0})$} in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
@@
This $g$ satisfies the fully coupled quadratic condition $(H_{q}^{\prime})$, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7. First, we note that the terminal condition in BSDE 6.37 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.39 in terms of the centered random variable $\eta$:
@@
$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi].$
@@
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies  $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
@@
Then the BSDE 6.39 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
@@
$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}dW_{s}.$
@@
$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\left\|U\right\|_{\mathbb{S}^{\infty}}\mathbb{E}\left[\int_{t}^{T}\Big|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\Big|ds\mid\mathcal{F}_{t}\right]$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|^{2}d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right].$ (6.41)
@@
Step 3. Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.43) is in fact more stringent than (6.45). Consequently, if we define $\rho>0$ by
@@
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
@@
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.40, the solution to the original BSDE 6.39 is obtained by the explicit transformation
@@
Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the conditional expectation, i.e., $g(s,Z_{s},\mathbb{E}[Z_{s}])$. In particular, the existence and uniqueness result applies to BSDEs of the type
@@
which will be useful for analyzing the driver (6.38) in this section.
While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$ as in hibon2017MFquadratic. In the present section, however, the application of interest involves drivers depending solely on $Z$ and its expectation, so the above results apply directly. We now present the main result of this section.
@@
First we will show the driver $f$ (6.38) satisfies the condition $(\mathrm{H}_{q}^{\prime})$. To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
@@
Together with the boundedness of $\theta$, we obtain, for some constants $c_{1},c_{2}>0$, for all $t\in[0,T]$
$\quad\leq c_{1}\,|Z_{t}^{1}-Z_{t}^{2}|+c_{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
which is a sufficient condition for $(\rm H_{q}^{\prime})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.37) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
@@ 7 Conclusion
In this paper we have developed a general framework for quadratic mean-field BSDEs whose generators exhibit quadratic growth in the control and its law. Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition. The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators.
We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms. In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition. As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.

### v2

@@ 6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.32 admits a unique adapted solution. Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.30 guarantees the existence and uniqueness of solution. Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.32) with the same formulation and the same $(\rm H_{0})$} condition in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
@@
This $g$ satisfies the fully coupled quadratic condition $(\rm H_{q}^{\prime})$}, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
First, we note that the terminal condition in BSDE 6.30 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.32 in terms of the centered random variable $\eta$:
@@
$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi]$
are both well-defined because of $(\rm H_{0})$}.
@@
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
@@
Then the BSDE 6.33 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
@@
$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}\!\cdot dW_{s}.$
@@
$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right]^{2}.$ (6.34)
@@
Step 3. Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.36) is in fact more stringent than the smallness condition (6.38). Consequently, if we define $\rho>0$ by
@@
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
@@
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.33, the solution to the original BSDE 6.32 is obtained by the explicit transformation
@@
Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the expectation $\mathbb{E}[Z]$ instead. In particular, the existence and uniqueness result applies to BSDEs of the type
@@
which will be useful for analyzing the driver (6.31) we picked earlier in the context of utility maximization in this section.
While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$. In the present section, however, the application of interest involves drivers depending solely on $Z$ and its expectation, so the above results apply directly. We now present the main result of this section.
@@
First we will show the driver $f$ (6.31) satisfies the condition $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. The verification of the first condition is straightforward due to the boundedness of $\theta$. To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
@@
Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\rm H_{q}^{\prime})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.30) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
@@ 7 Conclusion
In this paper we have developed a general framework for mean-field BSDEs whose generators have quadratic growth in the control and its law. Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition. The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators. We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms. In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition. As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.